The standard deviation of the market-index portfolio is 20%. Stock A has a beta of 2.50 and a residual standard deviation of 30%.

a. Calculate the total variance for an increase of 0.25 in its beta. (Do not round intermediate calculations. Round your answer to the nearest whole number.)

Total Variance =

b. Calculate the total variance for an increase of 7.75% in its residual standard deviation. (Do not round intermediate calculations. Round your answer to the nearest whole number.)

Total Variance =
Answer #1

a)
Total variance = (2.75)^2 * (0.20)^2 + (0.30)^2
= 7.5625 * 0.04 + 0.09
= 0.3025 + 0.09
= 0.3925

b)
Total variance = (2.50)^2 * (0.20)^2 + (0.3775)^2
= 6.25 * 0.04 + 0.1425
= 0.25 + 0.1425
= 0.3925